# Occupancy-SLAM: An Efficient and Robust Algorithm for Simultaneously Optimizing Robot Poses and Occupancy Map

### v2

@@ VIII-B2 Datasets
We perform comparisons using three real-world datasets. (1) The Newer College Dataset [60]: The first five sequences from the shorter experiment, collected with a handheld Ouster OS-1 LiDAR scanner at New College, Oxford. The environment includes lawns, buildings, a tunnel, and a garden. Ground truth is provided by a BLK360 LiDAR scanner to capture a detailed 3D map and then infer the ground truth of poses with centimeter-level accuracy. (2) KITTI Dataset [61]: Sequence 07, a demo dataset for HBA, collected with a Velodyne HDL-64E LiDAR scanner mounted on a car. Ground truth poses is provided by RTK-GPS/INS. (3) Arche Dataset [33]: A demo dataset for Voxgraph, collected using an Ouster OS1 LiDAR mounted on a hexacopter MAV in a disaster area. Ground truth positions are provided by an RTK-GNSS system.
The Newer College Dataset is used to evaluate high-precision performance in normal-scale environments, while the KITTI and Arche datasets are used to test performance in large environments with long trajectories.
Figure 16: Some local point cloud maps from the Arche dataset. The first row shows point cloud maps generated using odometry from ROVIO (also used for submap construction in Voxgraph), while the second row shows maps generated with our optimized poses using the same LiDAR scans. BALM2 fails to produce results when using the same odometry and scans as inputs in all these local environments.
@@ VIII-B3 Experiments on normal-scale environments
We evaluate the performance of our proposed method without submap joining in normal-scale environments.
First, we evaluate BALM2 and our method using the first five sequences of The Newer College Dataset, which encompass all scenarios within the dataset. As shown in Table IX, our method outperforms BALM2 across all metrics, except for the RMSE in Seq. 1, and significantly outperforms the odometry inputs from FAST-LIO2 in all metrics.
@@
Next, we test robustness in a challenging environment with noisy odometry input using the Arche dataset. This dataset, collected by a hexacopter MAV in an unstructured environment, is influenced by drone vibrations, flight speed, and environmental factors. Local point cloud maps built using odometry from ROVIO [62] (also used to construct submaps in Voxgraph) are shown in the first row of Fig. 16. To evaluate BALM2 and our method, we partition the dataset into several short sequences, each lasting 10–20 seconds. BALM2 fails in all the sequences except during MAV start-up and landing due to insufficient planar features for optimization, while our method performs well on all the sequences. The second row of Fig. 16 illustrates some of our results, demonstrating that our method is robust in 3D and does not rely on environmental assumptions. Additionally, the results confirm that our method achieves significantly higher pose accuracy than ROVIO.
@@ VIII-B4 Experiments on large-scale environments
We evaluate our method in large-scale environments with long trajectories using the KITTI and Arche datasets, comparing it with HBA and Voxgraph. For this experiment, we first build submaps by jointly optimizing poses and maps within submaps, then apply our submap joining algorithm to jointly optimize submap frame poses and the global occupancy map.
Table X summarizes the MAE and RMSE of absolute trajectory error, showing our method achieves the best results on both datasets. Fig. 17 illustrates that our approach can achieve the best global robot trajectories. Notably, our method significantly outperforms Voxgraph on the KITTI dataset and HBA on the Arche dataset. The relatively poor performance of Voxgraph on the KITTI dataset is due to its reliance on relative measurements from SDF-to-SDF registration, which requires sufficient overlapping submaps—a challenge in autonomous driving scenarios. In contrast, our submap joining algorithm jointly optimizes submap poses and the global occupancy map, avoiding this limitation. HBA underperforms on the Arche dataset due to its reliance on planar features for optimization, which is challenging in unstructured environments and during MAV motion. Odometry and point clouds from such scenarios make detecting sufficient planar features difficult, and the planarity assumption often fails in non-urban environments like disaster areas.
@@
Figure 17: Robot trajectory results of datasets in large-scale environments. (a) and (b) show the trajectories of ground truth, Voxgraph [33], HBA [49], and our method for KITTI dataset and Arche dataset.
@@ VIII-C Discussion
The experimental results in this section demonstrate that our proposed idea of jointly optimizing the robot pose and the occupancy map can also lead to better solutions for the robot poses and occupancy maps in 3D cases. However, several challenges remain in 3D scenarios. For instance, 3D point clouds from LiDAR scanners are often sparse, particularly in the vertical direction, which can lead to observability issues in the optimization problem. This sparsity also results in inhomogeneous observation information, complicating the accurate representation of the 3D environment in occupancy maps. Furthermore, the large dimensions of 3D maps pose significant computational challenges in large-scale SLAM, requiring more efficient solving methods.
To address these challenges, several potential solutions can be explored. First, adopting compact representations for 3D occupancy maps, such as octree structures similar to Octomap [8] and [21], can enhance efficiency. Second, combining local map and pose optimization with hierarchical optimization and submap joining methods can further reduce computational time. Lastly, using continuous representations for 3D occupancy maps enables more precise gradient calculations, which can better guide the optimization process.
@@ IX Conclusion
In this paper, we propose Occupancy-SLAM algorithm, which solves robot poses and occupancy map simultaneously. To enhance efficiency and robustness, we introduce a multi-resolution strategy. The first stage jointly optimizes poses and a low-resolution occupancy map to quickly achieve relatively accurate pose estimates, which are then used as the initial guess for the second stage. The second stage refines poses and a subset of the high-resolution map, focusing on critical boundary areas. Additionally, we extend this framework to an occupancy grid-based submap joining algorithm, addressing challenges in large-scale environments and long-term trajectories. Results from both simulated and real-world datasets demonstrate that our method achieves more accurate pose and map estimates than state-of-the-art approaches.
@@
In our future work, we will further explore problem formulation and solution techniques in the 3D case to develop more efficient and robust algorithms capable of addressing various challenges in 3D environments.
@@
The Jacobian $\mathbf{J}_{P}$ of function $F_{ij}^{Z}(\mathbf{x})$ in the observation term w.r.t. the robot poses $\mathbf{x}^{P}_{i}$ can be calculated by the chain rule
$\mathbf{J}_{P}=\frac{\partial F_{ij}^{Z}(\mathbf{x})}{\partial\mathbf{x}^{P}_{i}}=\frac{\partial F_{ij}^{Z}(\mathbf{x})}{\partial\mathbf{p}_{ij}}\cdot\frac{\partial\mathbf{p}_{ij}}{\partial\mathbf{x}^{P}_{i}}$ (28)
in which $\dfrac{\partial\mathbf{p}_{ij}}{\partial\mathbf{x}^{P}_{i}}$ can be calculated as
$\dfrac{\partial\mathbf{p}_{ij}}{\partial\mathbf{x}^{P}_{i}}=\left[\begin{array}[]{ll}\dfrac{\partial\mathbf{p}_{ij}}{\partial\mathbf{t}_{i}}&\dfrac{\partial\mathbf{p}_{ij}}{\partial\theta_{i}}\end{array}\right]=\dfrac{1}{s}\left[\begin{array}[]{ll}\mathbf{E}_{2}&\left(\mathbf{R}_{i}^{\prime}\right)^{\top}\mathbf{p}_{ij}\end{array}\right].$ (29)
$\mathbf{R}_{i}^{\prime}$ is the derivative of the rotation matrix $\mathbf{R}_{i}$ w.r.t. rotation angle $\theta_{i}$ and $\mathbf{E}_{2}$ means $2\times 2$ identity matrix.
@@
Here $\dfrac{\partial M(\mathbf{p}_{ij})}{\partial\mathbf{p}_{ij}}$ can be considered as the gradient of the occupancy map at point $\mathbf{p}_{ij}$, which can be approximated by the bilinear interpolation of the gradients of the occupancy at the four adjacent cell vertices $\mathbf{\nabla}M(\mathbf{m}_{wh}),\cdots,\mathbf{\nabla}M(\mathbf{m}_{(w+1)(h+1)})$ around $\mathbf{p}_{ij}$ as
$\dfrac{\partial M(\mathbf{p}_{ij})}{\partial\mathbf{p}_{ij}}=\left[\begin{aligned} a_{1}b_{1}\\ a_{0}b_{1}\\ a_{1}b_{0}\\ a_{0}b_{0}\\ \end{aligned}\right]^{\top}\left[\begin{aligned} &\mathbf{\nabla}M(\mathbf{m}_{wh})\\ &\mathbf{\nabla}M(\mathbf{m}_{(w+1)h})\\ &\mathbf{\nabla}M(\mathbf{m}_{w(h+1)})\\ &\mathbf{\nabla}M(\mathbf{m}_{(w+1)(h+1)})\end{aligned}\right]$ (31)
where the gradient of occupancy map $\mathbb{M}$ at all the cell vertices $\mathbf{\nabla}M$ can be easily calculated from $\mathbf{x}^{M}$ in the state. The bilinear interpolation used in (31) is similar to the method in (LABEL:eq_interp).
Here, we assume the robot poses $\mathbf{x}^{P}$ change slightly in each iteration, to reduce the computational complexity, the hit map $\mathbb{N}$ is considered as constant and recalculated using the current robot poses in each iteration. Thus, the derivative of $N(\mathbf{p}_{ij})$ is not calculated.
@@
Based on (LABEL:eq_interp), the Jacobian $\mathbf{J}_{M}$ of function $F_{ij}^{Z}(\mathbf{x})$ in the observation term w.r.t. the map part of state vector $\mathbf{x}^{M}$ can be calculated as
@@
The Jacobian $\mathbf{J}_{O}$ of function $F_{i}^{O}(\mathbf{x})$ in the odometry term (14) is the partial derivative w.r.t. the robot poses $\mathbf{x}^{P}$ since it is not related to the occupancy map in the state vector $\mathbf{x}$. Therefore, the Jacobian $\mathbf{J}_{O}$ can be calculated as
@@
The Jacobian $\mathbf{J}_{S}$ of function $F^{S}(\mathbf{x})$ in the smoothing term is the derivative of (15) w.r.t. cell vertices of occupancy map $\mathbf{x}^{M}$ due to it is not related to the robot poses $\mathbf{x}^{P}$ in the state vector $\mathbf{x}$. It should be mentioned that $F^{S}(\mathbf{x})$ is linear w.r.t. $\mathbf{x}^{M}$
@@
where the $(2c_{w}c_{h}+c_{w}+c_{h})\times((c_{w}+1)(c_{h}+1))$ coefficient matrix $\mathbf{A}$ is sparse and with nonzero elements $1$ or $-1$. An example of the coefficient matrix can be shown as
@@
Here $\mathbf{0}$ represents a zero vector with appropriate dimensions. Therefore, the Jacobian of the smoothing term can be calculated as
@@
For Jacobian $\mathbf{J}_{P}$, partial derivatives w.r.t. all the cell vertices are required for (31). However, not all vertices are included in the state vector in the second stage, which makes it challenging to compute the partial derivatives w.r.t. some cell vertices because their surrounding nodes are discarded. From a semantic perspective, the discarded cell vertices have the same occupancy state as the edge nodes, which is why they are excluded. Consequently, the gradient of these edge vertices is expected to be close to zero. Based on this reasoning, we set the partial derivatives w.r.t. all edge cell vertices to 00 when they need to be calculated using (31).
For Jacobian $\mathbf{J}_{S}$, it can also be calculated using the same idea as (36). In the second stage of our multi-resolution strategy, (36) is reformulated as

### v3

@@ VIII-B2 Datasets
We evaluate our method on three real-world datasets. (1) The Newer College Dataset [60]: The first five sequences from the shorter experiment, collected with a handheld Ouster OS-1 LiDAR scanner at New College, Oxford. The environment includes lawns, buildings, a tunnel, and a garden. Ground truth poses are derived from high-precision 3D maps constructed using a BLK360 LiDAR scanner, achieving centimeter-level accuracy. (2) KITTI Dataset [61]: Sequence 07, a demo dataset used in HBA, recorded with a vehicle-mounted Velodyne HDL-64E LiDAR. Ground truth poses are provided by RTK-GPS/INS. (3) Arche Dataset [33]: A demo dataset for Voxgraph, collected using an Ouster OS1 LiDAR mounted on a hexacopter MAV in a disaster area. Ground truth positions are provided by an RTK-GNSS system.
The Newer College Dataset is used to evaluate high-precision performance in normal-scale environments, while the KITTI and Arche datasets are used to assess performance in large environments with long trajectories.
Figure 16: Some local point cloud maps from the Arche dataset. The first row shows point cloud maps generated using odometry from ROVIO, which is also used for submap construction in Voxgraph. The second row shows maps generated with our optimized poses using the same LiDAR scans. BALM2 fails to produce results in all these local environments when given the same odometry and scan inputs.
@@ VIII-B3 Experiments on normal-scale environments
We evaluate the performance of our method without submap joining in normal-scale environments.
We first compare FAST-LIO2, BALM2, and our method on the first five sequences of The Newer College Dataset, which encompass all scenarios within the dataset. As shown in Table IX, our method outperforms BALM2 across all metrics except RMSE in Seq. 1, and significantly outperforms the odometry inputs from FAST-LIO2 in all metrics.
@@
We then assess robustness under challenging conditions using the Arche dataset, collected by a hexacopter MAV in an unstructured environment affected by vibrations, flight dynamics, and external disturbances. Local point cloud maps built using odometry from ROVIO [62] (also used to construct submaps in Voxgraph) are shown in the top row of Fig. 16. For evaluation, we partition the dataset into several short sequences, each lasting 10–20 seconds. BALM2 fails on all but the initial and final segments due to a lack of planar features of adequate quality for optimization, whereas our method performs reliably across all sequences. The bottom row of Fig. 16 illustrates some of our results, demonstrating that our method is robust in 3D and does not rely on environmental assumptions. Additionally, the results confirm that our method achieves significantly higher pose accuracy than ROVIO.
@@ VIII-B4 Experiments on large-scale environments
We evaluate our method in large-scale environments with long trajectories using the KITTI and Arche datasets, comparing it with HBA and Voxgraph. For this experiment, we first build submaps by jointly optimizing poses and maps within each submap, then apply our submap joining algorithm to jointly optimize submap frame poses and the global occupancy map.
Table X summarizes the absolute trajectory error (in MAE and RMSE), showing that our method achieves the best performance on both datasets. Fig. 17 illustrates that our approach achieves the most accurate global trajectories. Notably, our method significantly outperforms Voxgraph on the KITTI dataset and HBA on the Arche dataset. Voxgraph performs poorly on KITTI due to its reliance on relative measurements from SDF-to-SDF registration, which requires substantial overlap between submaps—often lacking in autonomous driving scenarios. In contrast, our submap joining algorithm jointly optimizes submap poses and the global occupancy map, avoiding this limitation. HBA underperforms on the Arche dataset due to its dependence on planar features, which are difficult to extract in unstructured environments and during MAV motion. In such scenarios, odometry and point cloud quality degrade, and the planarity assumption often breaks down in non-urban environments such as disaster areas.
@@
Figure 17: Robot trajectory results in large-scale environments. (a) and (b) show the trajectories of ground truth, Voxgraph [33], HBA [49], and our method on the KITTI and Arche datasets, respectively.
@@ VIII-C Discussion
The experimental results in this section demonstrate that our proposed idea of jointly optimizing the robot pose and the occupancy map also yields improved solutions in 3D scenarios. However, several challenges remain. For instance, 3D LiDAR point clouds are often sparse, particularly in the vertical direction, which can cause observability issues in the optimization. This sparsity also results in inhomogeneous observation density, complicating the accurate representation of the 3D environment in occupancy maps. Furthermore, the high dimensionality of 3D maps introduces significant computational burdens in large-scale SLAM, requiring more efficient solving methods.
To address these challenges, several potential solutions can be explored. First, adopting compact representations such as octree structures (e.g., OctoMap [8] and [21]) can enhance efficiency. Second, combining local map and pose optimization with hierarchical optimization and submap joining methods can further reduce computational cost. Finally, leveraging continuous representations of 3D occupancy maps enables more precise gradient computation, thereby providing better guidance for optimization.
@@ IX Conclusion
In this paper, we propose Occupancy-SLAM algorithm, which solves robot poses and occupancy map simultaneously. To enhance efficiency and robustness, we introduce a multi-resolution strategy. The first stage jointly optimizes poses and a low-resolution occupancy map to quickly achieve relatively accurate pose estimates, which then serves as the initial guess for the second stage. The second stage refines poses and a subset of the high-resolution map, focusing on critical boundary areas. We further extend this framework to an occupancy grid-based submap joining algorithm, addressing challenges in large-scale environments and long-term trajectories. Results from both simulated and real-world datasets demonstrate that our method achieves more accurate pose and map estimates than state-of-the-art approaches.
@@
In future work, we aim to further explore problem formulation and solution techniques for 3D environments, with the goal of developing more efficient and robust algorithms capable of addressing various challenges of 3D SLAM.
@@
The Jacobian of $F_{ij}^{Z}(\mathbf{x})$ in the observation term w.r.t. the robot poses $\mathbf{x}^{P}_{i}$ can be calculated using the chain rule
$\mathbf{J}_{P}=\frac{\partial F_{ij}^{Z}(\mathbf{x})}{\partial\mathbf{x}^{P}_{i}}=\frac{\partial F_{ij}^{Z}(\mathbf{x})}{\partial\mathbf{p}_{ij}}\cdot\frac{\partial\mathbf{p}_{ij}}{\partial\mathbf{x}^{P}_{i}}.$ (28)
Here, $\dfrac{\partial\mathbf{p}_{ij}}{\partial\mathbf{x}^{P}_{i}}$ is calculated as
$\dfrac{\partial\mathbf{p}_{ij}}{\partial\mathbf{x}^{P}_{i}}=\left[\begin{array}[]{ll}\dfrac{\partial\mathbf{p}_{ij}}{\partial\mathbf{t}_{i}}&\dfrac{\partial\mathbf{p}_{ij}}{\partial\theta_{i}}\end{array}\right]=\dfrac{1}{s}\left[\begin{array}[]{ll}\mathbf{E}_{2}&\left(\mathbf{R}_{i}^{\prime}\right)^{\top}\mathbf{p}_{ij}\end{array}\right],$ (29)
where $\mathbf{R}_{i}^{\prime}$ is the derivative of the rotation matrix $\mathbf{R}_{i}$ w.r.t. rotation angle $\theta_{i}$, and $\mathbf{E}_{2}$ means $2\times 2$ identity matrix.
@@
Here $\dfrac{\partial M(\mathbf{p}_{ij})}{\partial\mathbf{p}_{ij}}$ can be considered as the gradient of the occupancy map at point $\mathbf{p}_{ij}$, which can be approximated by bilinear interpolation of the gradients of the occupancy at the four adjacent cell vertices $\mathbf{\nabla}M(\mathbf{m}_{wh}),\cdots,\mathbf{\nabla}M(\mathbf{m}_{(w+1)(h+1)})$ around $\mathbf{p}_{ij}$ as
$\dfrac{\partial M(\mathbf{p}_{ij})}{\partial\mathbf{p}_{ij}}=\left[\begin{aligned} a_{1}b_{1}\\ a_{0}b_{1}\\ a_{1}b_{0}\\ a_{0}b_{0}\\ \end{aligned}\right]^{\top}\left[\begin{aligned} &\mathbf{\nabla}M(\mathbf{m}_{wh})\\ &\mathbf{\nabla}M(\mathbf{m}_{(w+1)h})\\ &\mathbf{\nabla}M(\mathbf{m}_{w(h+1)})\\ &\mathbf{\nabla}M(\mathbf{m}_{(w+1)(h+1)})\end{aligned}\right].$ (31)
The gradient $\nabla M(\cdot)$ at all cell vertices can be easily calculated from $\mathbf{x}^{M}$ in the state. The bilinear interpolation used in (31) is similar to the method in (LABEL:eq_interp).
We assume the robot poses $\mathbf{x}^{P}$ change slightly in each iteration. To reduce the computational complexity, the hit map $\mathbb{N}$ is considered as constant during each iteration and is recalculated based on the current robot poses. The derivative of $N(\mathbf{p}_{ij})$ is therefore omitted from the Jacobian computation.
@@
Based on (LABEL:eq_interp), the Jacobian of $F_{ij}^{Z}(\mathbf{x})$ in the observation term w.r.t. the map part in state vector $\mathbf{x}^{M}$ can be calculated as
@@
The Jacobian of $F_{i}^{O}(\mathbf{x})$ in the odometry term (14) is the partial derivative w.r.t. the robot poses $\mathbf{x}^{P}$ since it is not related to $\mathbf{x}^{M}$. Therefore, the Jacobian $\mathbf{J}_{O}$ can be calculated as
@@
The Jacobian $\mathbf{J}_{S}$ is the derivative of (15) w.r.t. $\mathbf{x}^{M}$ due to it is not related to the robot poses $\mathbf{x}^{P}$. It should be mentioned that $F^{S}(\mathbf{x})$ is linear w.r.t. $\mathbf{x}^{M}$, i.e.,
@@
where the coefficient matrix $\mathbf{A}$ is sparse and with nonzero elements $1$ or $-1$. An example of the coefficient matrix can be shown as
@@
Here $\mathbf{0}$ represents a zero vector with appropriate dimensions. The Jacobian of the smoothing term can be calculated as
@@
For Jacobian $\mathbf{J}_{P}$, partial derivatives w.r.t. all the cell vertices are required for (31). However, not all vertices are included in the state vector in the second stage, which makes it challenging to compute the partial derivatives w.r.t. some cell vertices because their surrounding nodes are discarded. From a semantic perspective, the discarded cell vertices share the same occupancy state as the edge vertices, which justifies their exclusion. Consequently, the gradients at these edge vertices are assumed to be negligible. Based on this assumption, the partial derivatives w.r.t. all edge cell vertices are set to zero when computing (31).
Jacobian $\mathbf{J}_{S}$ can be calculated using the same idea as (36). In the second stage of our multi-resolution strategy, (36) is reformulated as